Allison:

A Ferris wheel has a diameter of 225 feet. If a passenger gets in a car and travels 38 feet when the wheel stops to let more passengers on, find the angle of rotation to the nearest

dude:

To the nearest degree? Draw it :)  In this case, they give you arc length, you are solving for the angle \(s=2\pi r (\dfrac{\theta}{360})\) s = arc length = 38 radius is diameter / 2 which in this case is \(\dfrac{225}{2}\) We solve for greek letter \(38=2\pi (112.5)(\dfrac{\theta}{360})\)

Allison:

112.5 is r

dude:

Yes

Allison:

Greek letter is 38?

dude:

We are solving for greek letter this time \(38=2\pi (112.5)(\dfrac{\theta}{360})\)

Allison:

Oh

Allison:

What do I do with the 38?

dude:

Its equal to all of that \(\color{red}{38}=2\pi (112.5)(\dfrac{\theta}{360})\)

Allison:

o.0

Allison:

So I divide instead?

dude:

Divide by 2 pi (112.5) and multiply 360 by 38 \(\dfrac{38\cdot 360}{2\pi (112.5)}=\theta\)

Allison:

I see what you're saying

Allison:

38/ 2pi(112.5)?

dude:

Yes and multiply by 360

Allison:

6715.15 ?

dude:

Hmm no

dude:

Make sure you put everything correctly s:

Allison:

19.35

dude:

Yesss Does it ask you to round to the nearest what though

Allison:

!!!

dude:

It got cut off

Allison:

Omg it's all wrong because I didn't put measurements

Allison:

Oh you're joking

Allison:

Omg

dude:

Boiii